MATLAB Question: Evaluate the polynomial expression 3x^3+4x^2+2x?2 at x=4,6,and x=8

MATLAB Question:

Evaluate the polynomial expression 3x^3+4x^2+2x?2 at x=4,6,and x=8
Answer #1

x = 4

= 3 x^3+4 x^2+2 x^2
= 3 *4^3 + 4 *4^2 + 2 *4^2
= 3 * 64 + 4 * 16 + 2 * 16
= 192 + 64 + 16
= 272


x = 6

= 3 x^3+4 x^2+2 x^2
= 3 * 6^3 + 4 *6^2 + 2 *6^2
= 3 * 216 + 4 * 36 + 2 * 36
= 648 + 144 + 72
= 864

x = 8

= 3 x^3+4 x^2+2 x^2
= 3 * 8^3 + 4 *8^2 + 2 *8^2
= 3 * 256 + 4 * 64 + 2 * 64
= 768 + 256 + 128
= 1152